OpenStudy (anonymous):

4a+7b=-5 8a+4c=36 6b+4c=-14

OpenStudy (anonymous):

what are trying to do here? just solve the system of equations?

OpenStudy (anonymous):

nvm that can you classify independent dependent and incosistent systsems

OpenStudy (anonymous):

2x-y=3 6x-3y=9

OpenStudy (anonymous):

nvm? what does that mean

OpenStudy (anonymous):

never mind can you help me with the one i just asked im doing a timed test!!

OpenStudy (anonymous):

this system just cancels out when you try to solve it

OpenStudy (anonymous):

its inconsistent

OpenStudy (anonymous):

thanksss sooo much are you busy? can you continue to help me

OpenStudy (anonymous):

I'll try to

OpenStudy (anonymous):

okay solve the system by elimination

OpenStudy (anonymous):

4a+4b-c=-6 8a+6b+c=-5 6a-4b-6c=-22

OpenStudy (anonymous):

I'm sorry I don't remember how to use elimination when you have 3 of them

OpenStudy (anonymous):

its okay how bout this one x+y=-4 x-y+6

OpenStudy (anonymous):

is that supposed to be =6

OpenStudy (anonymous):

instead of +6

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

if so the y's cancel out....so x=1 and y=-5

OpenStudy (anonymous):

on the 2nd equation you have +6 and I think you mean =6

OpenStudy (anonymous):

no! you are confusing me

OpenStudy (anonymous):

this is what you typed a few minutes ago......."its okay how bout this one x+y=-4 x-y+6"